OpenStudy (anonymous):

What is the equation in point slope form of the line passing through (0,2) and (1,5) Y-5=3(x-1) Y-2= -3(x-0) Y+2=3(x+1) Y+5=-3(x+1)

OpenStudy (jdoe0001):

can you find the slope of it?

OpenStudy (anonymous):

find the slope: (5-2)/(1-0)=3/1=3 Then plug it into the point slope formula: y-y1=m(x-x1) So using the first point: y-2=3(x-0)